$SPAD/src/input c02agf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
a:Matrix SF:=
 [[1.0 ,2.0 ,3.0 ,4.0 ,5.0 ,6.0 ]]
n:=ncols(a)-1
scale:=true
result:=c02agf(a,n,scale,-1)

———-

References^{[1] nothing}

4